OpenStudy (anonymous):

(sin^-1(x))/sqrt (1-x^(2))

OpenStudy (jdoe0001):

is there an equal sign somewhere?

OpenStudy (amoodarya):

\[\int\limits \frac{ \arcsin x }{ \sqrt{1-x^2} }dx=\\ ?\]

OpenStudy (amoodarya):

if yes use substitution u = arcsin x

OpenStudy (amoodarya):

\[\int\limits \frac{ \arcsin x }{ \sqrt{1-x^2} }dx=\\ u=\arcsin x \\du =\frac{ 1 }{ \sqrt{1-x^2}}dx\\\]

OpenStudy (amoodarya):

its turn to \[\int\limits u du = \frac{ u^2 }{2 }+c = \frac{ (arcsinx)^2 }{2 }+c\]